OpenStudy (anonymous):

1. Plot the points A(-3,0) B(3,2) C(5,-4) 2. Find the slope of AB, BC, and AC (Show work)

ganeshie8 (ganeshie8):

do you have a graph sheet or something similar to plot the points ?

OpenStudy (anonymous):

Yes @ganeshie8

ganeshie8 (ganeshie8):

have you plotted them ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Oh I meant yes to plotting and no to finding the slope

ganeshie8 (ganeshie8):

good, use the slope formula for finding the slope

OpenStudy (anonymous):

Thank you

ganeshie8 (ganeshie8):

A(-3,0) B(3,2) slope of AB would be \[\dfrac{2-0}{3-(-3)}\]

ganeshie8 (ganeshie8):

\[\dfrac{2}{6}\] which is same as \[\dfrac{1}{3}\]

OpenStudy (anonymous):

Cause 2/6 simplified is 1/3 right?

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

in your plot you will see that the line rises UP by 1 unit when you run 3 units to the RIGHT

ganeshie8 (ganeshie8):

here it is http://gyazo.com/504b0d1af1680171d5219406d2442e84